# A Survey on Bundle Recommendation: Methods, Applications, and Challenges

### v1

@@ 5.3. Comparison and Experiments
We conduct reproducibility experiments on bundle recommendation. For discriminative bundle recommendation, we reproduce 6 mainstream graph-based models and 4 CF-based models, and compare their performance. For generative bundle recommendation, we focus on trip generation and recommendation, a representative generative scenario with sufficient resources to support empirical experiments.
@@ 5.3.1. Discriminative BR
Datasets and Evaluation Metrics.
We use the following three real-world datasets for evaluation, with their statistics shown in Table 10.
Table 10. Dataset Statistics of Discriminative BR
Dataset #U #I #B #U-I #U-B #Avg.I/B
Youshu 8,039 32,770 4,771 138,515 51,377 37.03
NetEase 18,528 123,628 22,864 1,128,065 302,303 77.80
iFashion 53,897 42,563 27,694 2,290,645 1,679,708 3.86
Two widely used metrics (Chang et al., 2020; He et al., 2019b; Zhao et al., 2022), Recall@K and NDCG@K, are adopted to evaluate the top-K recommendation performance, where $K\in\{20,40\}$.
Implementation Details.
We reproduce 6 graph-based models, CrossCBR (Ma et al., 2022), MultiCBR (Ma et al., 2024a), BundleGT (Wei et al., 2023), MIDGN (Zhao et al., 2022), BGCN (Chang et al., 2020), and HyperMBR (Ke et al., 2023a), 4 CF-based models, LightGCN (He et al., 2020a), NGCF (Wang et al., 2019), ItemKNN (Deshpande and Karypis, 2004), MFBPR(Rendle et al., 2009), on the Youshu, NetEase and iFashion datasets. Among these, MultiCBR, BundleGT, MIDGN, BGCN and HyperMBR are reproduced on the three datasets using five different random seeds (123, 124, 125, 126, 127) and take the average as the result. For all methods, we follow the optimal parameters provided in the cited papers. We adopt BPR loss, employ Adam to optimize parameters and set the embedding size as 64. All model training and testing are performed on NVIDIA A40.
Performance Comparison.
The performance comparison of six reproduced models on three datasets is presented in Table 11. Based on these results, we have the following observations:
@@ 5.3.2. Generative BR
Datasets and Evaluation Metrics.
In the TripRec task, real-world POI check-in datasets from Flickr and Foursquare are often widely used for experimental evaluations (Gao et al., 2022; Chen et al., 2016; Zhao et al., 2021), thus we utilize these for experiments. Eight cities are selected from the datasets, i.e., Edinburgh, Glasgow, Osaka, Toronto, Budapest, Vienna, Melbourne, and Tokyo. The basic dataset statistics are shown in Table 12, where the name of each city has a corresponding abbreviation.
Table 12. Dataset Statistics of Generative BR
City Name
Edinburgh
Glasgow
Osaka
Toronto
Budapest
Vienna
Melbourne
Tokyo
#User 1, 454 601 450 1, 395 935 1, 155 1, 000 200
#POI Visits 33, 944 11, 434 7, 747 39, 419 18, 513 34, 515 23, 995 8, 721
#Trajectory 5,028 2,227 1,115 6,057 2,361 3,193 5,106 6,414
The precision, recall, and F1 score are commonly used in state-of-the-art TripRec methods (Wang et al., 2023a; Lim et al., 2015; Zhao et al., 2021). Considering that F1 score is calculated based on precision and recall, therefore the well-established metric of F1 score is chosen to evaluate performance.
Implementation Details.
We choose and reproduce 12 models: (i) Traditional TripRec: RANDOM (Chen et al., 2016), PersTour (Lim et al., 2015), RANK+Markov (Chen et al., 2016), POIRANK (Chen et al., 2016), and TRAR (Gu et al., 2020). (ii) Neural TripRec: C-ILP (He et al., 2019a), C-ALNS (He et al., 2019a), CATHI (Zhou et al., 2019), DeepTrip (Gao et al., 2023b), PT-VPMF (Zhao et al., 2021), CTLTR (Zhou et al., 2022), and Query2Trip (Wang et al., 2023a). The The hyperparameter settings of our model are as follows: we set POI/trajectory embedding dimension as 256, batch size as 16, time embedding dimension as 32, hidden size as 128, learning rate as 0.5 (Trm), 0.01 (DAL), and 0.001 (DCL) respectively, the number of encoder layer as 5, and the number of attention heads as 8. We run model five times to average the results for experimental analysis.
Performance Comparison.
From the results in Table 13, we can make the following observations:
@@ 6. Challenges and Trends
While existing works have established a solid foundation for bundle recommendation research, there are still some further opportunities. In this section, we discuss some the main challenges and future trends.
@@ 6.1. Representation Learning in Non-Euclidean Spaces
Additionally, studies (Sun et al., 2024c, 2023c) have began to explore the theories and methods of graph learning in Riemannian space, a type of non-Euclidean space that can encompass different curvatures (positive, zero, and negative curvatures). CSincere (Sun et al., 2024c) discovered that the users and items differ in terms of both $\delta$-hyperbolicity and degree distribution, and believes that it is more rational to model the users and items in two different spaces. Riemannian geometry provides the notion of curvature to distinguish the structure pattern between different spaces. In bundle recommendation, as shown in Figure 8 and Table 14, the degree distributions of users, bundles, and items are different, and the $\delta$ values of the three types of interaction graphs are not the same. Therefore, modeling different entities in different spaces is reasonable and feasible. In addition, previous studies, whether in Euclidean or hyperbolic space, have tried to learn embeddings in a fixed curvature space. However, user interactions will continue to generate, requiring a dynamically changing curvature space to represent learning for users, items, and bundles. In summary, future research directions can be explored in two aspects: First, researchers can enrich the content of the datasets, supplement information for the existing mainstream datasets used in bundle recommendation research, such as adding community attributes and temporal dynamics, or build such datasets to lay the foundation for subsequent representation learning in adaptive Riemannian space. Second, researchers can consider using adaptive Riemannian space for representation learning. Moreover, considering that the curvature of the graph may change over time, incorporating the variation of graph curvature into the bundle recommendation model can make it more flexible to adapt to environmental changes, improving the model’s adaptability and accuracy.
[…]
Three entities degree distribution of three datasets
Figure 8. User, Bundle, Item Degree Distribution of Three Datasets
Table 14. $\delta$-hyperbolicity of different types of interaction graphs in NetEase, Youshu and iFashion
Datasets NetEase Youshu iFashion
U-B U-I B-I U-B U-I B-I U-B U-I B-I
${\delta}_{avg}$ 0.2846 0.3290 0.3406 0.3227 0.3341 0.3363 0.3020 0.3248 0.4135
@@ 6.2. Data Sparsity and Cold-Start Issues of Discriminative BR
Data sparsity and cold-start problems remain major challenges. The problem of skewed distribution of bundle interaction is prevalent by analyzing three mainstream datasets (Jeon et al., 2024), i.e., a small number of bundles get the vast majority of interactions, while a large number of bundles have little or no interactions at all. When new bundles are added to the system, it is difficult for the system to recommend them effectively due to the lack of historical interaction data. Recent research (Jeon et al., 2024) has proposed a solution for cold-start bundle recommendation, which solves the problem of highly skewed distribution of bundle interactions by taking into account the popularity of the bundles, and relying more on the user-item view, which provides richer information, for less popular bundles. To alleviate the data sparsity and cold-start problem in bundle recommendation, a promising future direction is to introduce multimodal information (e.g., images, descriptions of items or bundles, etc. ) to model user preferences harnessing LLMs. There have been many studies (Ma et al., 2024b; Liu et al., 2024; Kim et al., 2024; Ma et al., 2024c) that have began to introduce multimodal information and attempt to utilize LLMs to boost recommendation performance.
@@ 6.3. User Intents and Dynamic Bundle Generation of Generative BR
There have been studies (Sun et al., 2024b, a, 2023a; Li et al., 2023d) focusing on exploring the bundle generation that reflect user intents. The remarkable advancement of Large Language Models (LLMs) has led to widespread adoption for more effective recommendation (Harte et al., 2023; Dai et al., 2023; He et al., 2023). Notably, several studies (Dai et al., 2023; He et al., 2023) highlight ChatGPT’s potential to mitigate the cold start issue and provide explainable recommendations. Furthermore, studies (Sun et al., 2024b, 2023a) leverage the advanced reasoning capability of LLMs to infer user intent from user sessions through adaptive contextual learning paradigm and generate personalized bundles based on these intents. Inferring user intents not only helps to generate bundles that better meet the actual needs of users but is also key to enhancing the transparency of recommendation systems and user trust. Future directions can further explore how to more accurately capture and interpret user intents. On one hand, researchers can investigate how to design better contextual examples and prompts to enhance the interpretability of LLMs. On the other hand, future work could consider combining various types of data such as images and videos to provide richer explanations and recommendations. Moreover, considering that user preferences and intents change over time, the dynamic capture of user intent and the implementation of dynamic bundle generation become particularly important. The work (Yang et al., 2024b) proposes plans to explore dynamic bundle generation methods. Therefore, future research trends can focus on developing dynamic bundle generation methods that adapt to changes in user intent, and how to provide more interpretable recommendations through these methods.
@@ 6.4. Responsible Bundle Recommender System
Since a bundle is composed of several items, the bias issues in item recommendation can amplify issues such as bias and unfairness in bundle recommendation. We conduct an empirical study to analyze popularity bias on three datasets (Youshu, NetEase, iFashion). As shown in Figure 9(b), we find that the distribution of interactions is noticeably imbalanced, with some items receiving inadequate exposure, which exacerbates the unfairness among items and suggests the presence of bias in bundle recommendation. In addition, as discussed in the Section 6.2, these mainstream datasets generally suffer from a skewed distribution of user-bundle interactions, leading to a long-tail phenomenon where popular or frequently interacted bundles are recommended while neglecting less popular bundles, even though they better meet user needs. This popularity bias reduces fairness in recommendations an further marginalizes less popular bundles, exacerbating data imbalance.
(a) Bundle Interactions by Group (Item Interaction)
(b) Average Item Interactions by Group (Bundle Interaction)
Figure 9. Popularity bias analysis of three datasets (NetEase, Youshu, iFashion)
[…]
Popularity bias analysis of three datasets
@@ Acknowledgements.
This work is partially supported by NSFC, China (No.62276196) and the research grant (RGPIN-2020-07157 ) from the Natural Science and Engineering Research Council (NSERC) of Canada.

### v2

@@ 5.3. Comparison and Experiments
We conduct reproducibility experiments on bundle recommendation. For discriminative bundle recommendation, we reproduce 6 mainstream graph-based models and 4 CF-based models and compare their performance. For generative bundle recommendation, we focus on trip generation and recommendation, a representative scenario with sufficient resources for empirical experiments.
@@ 5.3.1. Discriminative BR
Experimental Details.
We use three real-world public datasets for evaluation, i.e., Youshu, NetEase, and iFashion. Two widely used metrics (Chang et al., 2020; He et al., 2019b; Zhao et al., 2022), Recall@K and NDCG@K, are adopted to evaluate the top-K recommendation performance, where $K\in\{20,40\}$. We reproduce 6 graph-based models, CrossCBR (Ma et al., 2022), MultiCBR (Ma et al., 2024a), BundleGT (Wei et al., 2023), MIDGN (Zhao et al., 2022), BGCN (Chang et al., 2020), and HyperMBR (Ke et al., 2023a), 4 CF-based models, LightGCN (He et al., 2020a), NGCF (Wang et al., 2019), ItemKNN (Deshpande and Karypis, 2004), MFBPR(Rendle et al., 2009), on the Youshu, NetEase and iFashion datasets. Among these, MultiCBR, BundleGT, MIDGN, BGCN and HyperMBR are reproduced on the three datasets using five different random seeds (123, 124, 125, 126, 127) and take the average as the result. For all methods, we follow the optimal parameters provided in the cited papers. We adopt BPR loss, employ Adam to optimize parameters and set the embedding size as 64. All model training and testing are performed on NVIDIA A40.
@@ 5.3.2. Generative BR
Experimental Details.
In the TripRec task, real-world POI check-in datasets from Flickr and Foursquare are often widely used for experimental evaluations (Gao et al., 2022; Chen et al., 2016; Zhao et al., 2021), thus we utilize these for experiments. Eight cities are selected from the datasets, i.e., Edinburgh, Glasgow, Osaka, Toronto, Budapest, Vienna, Melbourne, and Tokyo. The precision, recall, and F1 score are commonly used in state-of-the-art TripRec methods (Wang et al., 2023b; Lim et al., 2015; Zhao et al., 2021). Considering that F1 score is calculated based on precision and recall, therefore the well-established metric of F1 score is chosen to evaluate performance. We choose and reproduce 12 models: (i) Traditional TripRec: RANDOM (Chen et al., 2016), PersTour (Lim et al., 2015), RANK+Markov (Chen et al., 2016), POIRANK (Chen et al., 2016), and TRAR (Gu et al., 2020). (ii) Neural TripRec: C-ILP (He et al., 2019a), C-ALNS (He et al., 2019a), CATHI (Zhou et al., 2019), DeepTrip (Gao et al., 2023c), PT-VPMF (Zhao et al., 2021), CTLTR (Zhou et al., 2022), and Query2Trip (Wang et al., 2023b). The hyperparameter settings of our model are as follows: we set POI/trajectory embedding dimension as 256, batch size as 16, time embedding dimension as 32, hidden size as 128, learning rate as 0.5 (Trm), 0.01 (DAL), and 0.001 (DCL) respectively, the number of encoder layer as 5, and the number of attention heads as 8. We run model five times to average the results for experimental analysis.
@@ 6. Challenges and Trends
While existing works have established a solid foundation for bundle recommendation research, further opportunities remain. In this section, we discuss the main challenges and future trends.
@@ 6.1. Representation Learning in Non-Euclidean Spaces
Additionally, studies (Sun et al., 2024a, 2023a) have began to explore the theories and methods of graph learning in Riemannian space, a type of non-Euclidean space that can encompass different curvatures (positive, zero, and negative curvatures). CSincere (Sun et al., 2024a) discovered that the users and items differ in terms of both $\delta$-hyperbolicity and degree distribution, and believes that it is more rational to model the users and items in two different spaces. Riemannian geometry provides the notion of curvature to distinguish the structure pattern between different spaces. In bundle recommendation, the degree distributions of users, bundles, and items are different, and the $\delta$ values of the three types of interaction graphs are not the same. Therefore, modeling different entities in different spaces is reasonable and feasible. In addition, previous studies, whether in Euclidean or hyperbolic space, have tried to learn embeddings in a fixed curvature space. However, user interactions will continue to generate, requiring a dynamically changing curvature space to represent learning for users, items, and bundles. In summary, future research directions can be explored in two aspects: First, researchers can enrich the content of the datasets, supplement information for the existing mainstream datasets used in bundle recommendation research, such as adding community attributes and temporal dynamics, or build such datasets to lay the foundation for subsequent representation learning in adaptive Riemannian space. Second, researchers can consider using adaptive Riemannian space for representation learning. Moreover, considering that the curvature of the graph may change over time, incorporating the variation of graph curvature into the bundle recommendation model can make it more flexible to adapt to environmental changes, improving the model’s adaptability and accuracy.
@@ 6.2. Data Sparsity and Cold-Start Issues of Discriminative BR
Data sparsity and cold-start problems remain major challenges. The problem of skewed distribution of bundle interaction is prevalent by analyzing three mainstream datasets (Jeon et al., 2024; Li et al., 2025), i.e., a small number of bundles get the vast majority of interactions, while a large number of bundles have little or no interactions at all. When new bundles are added to the system, it is difficult for the system to recommend them effectively due to the lack of historical interaction data. Recent research (Jeon et al., 2024) has proposed a solution for cold-start bundle recommendation, which solves the problem of highly skewed distribution of bundle interactions by taking into account the popularity of the bundles, and relying more on the user-item view, which provides richer information, for less popular bundles. To alleviate the data sparsity and cold-start problem in bundle recommendation, a promising future direction is to introduce multimodal information (e.g., images, descriptions of items or bundles, etc. ) to model user preferences harnessing LLMs. There have been many studies (Ma et al., 2024b; Liu et al., 2024b; Kim et al., 2024; Ma et al., 2024c) that have began to introduce multimodal information and attempt to utilize LLMs to boost recommendation performance.
@@ 6.3. User Intents and Dynamic Bundle Generation of Generative BR
There have been studies (Sun et al., 2024c, b; Li et al., 2023c) focusing on exploring the bundle generation that reflect user intents. The remarkable advancement of Large Language Models (LLMs) has led to widespread adoption for more effective recommendation (Harte et al., 2023; Dai et al., 2023; He et al., 2023). Notably, several studies (Dai et al., 2023; He et al., 2023) highlight ChatGPT’s potential to mitigate the cold start issue and provide explainable recommendations. Furthermore, studies (Sun et al., 2024c) leverage the advanced reasoning capability of LLMs to infer user intent from user sessions through adaptive contextual learning paradigm and generate personalized bundles based on these intents. Inferring user intents not only helps to generate bundles that better meet the actual needs of users but is also key to enhancing the transparency of recommendation systems and user trust. Future directions can further explore how to more accurately capture and interpret user intents. On one hand, researchers can investigate how to design better contextual examples and prompts to enhance the interpretability of LLMs. On the other hand, future work could consider combining various types of data such as images and videos to provide richer explanations and recommendations. Moreover, considering that user preferences and intents change over time, the dynamic capture of user intent and the implementation of dynamic bundle generation become particularly important. The work (Yang et al., 2024b) proposes plans to explore dynamic bundle generation methods. Therefore, future research trends can focus on developing dynamic bundle generation methods that adapt to changes in user intent, and how to provide more interpretable recommendations through these methods.
@@ 6.4. Responsible Bundle Recommender System
Since a bundle is composed of several items, the bias issues in item recommendation can amplify issues such as bias and unfairness in bundle recommendation. We conduct an empirical study to analyze popularity bias on three datasets (Youshu, NetEase, iFashion). We find that the distribution of interactions is noticeably imbalanced, with some items receiving inadequate exposure, which exacerbates the unfairness among items and suggests the presence of bias in bundle recommendation. In addition, as discussed in the Section 6.2, these mainstream datasets generally suffer from a skewed distribution of user-bundle interactions, leading to a long-tail phenomenon where popular or frequently interacted bundles are recommended while neglecting less popular bundles, even though they better meet user needs. This popularity bias reduces fairness in recommendations and further marginalizes less popular bundles, exacerbating data imbalance.
@@
6.5. Large Language Models for Bundle Recommendation
In recommendation system, there has been a clear shift toward generative paradigms, where retrieval and ranking are reformulated as sequence modeling tasks (Rajput et al., 2023; Zhai et al., 2024). These advances highlight the potential of generative modeling to improve scalability, alleviate cold-start issues, and better capture sequential dependencies. Inspired by this trajectory, researchers have begun to investigate how such ideas can be tailored to bundle recommendation. Within bundle recommendation, a first line of work has leveraged multimodal foundation models to capture richer semantic information for bundle construction (Ma et al., 2024c; Nguyen et al., 2025; Ma et al., 2024b). By integrating features from text, images, audio, and user interaction data, these approaches uncover collaborative relations among items that extend beyond co-occurrence statistics. For example, methods such as RaMen (Nguyen et al., 2025) and CLHE (Ma et al., 2024c) employ multimodal encoders and contrastive learning to unify heterogeneous signals, thereby enhancing bundle representation learning and alleviating data sparsity. This demonstrates the importance of multimodal feature extraction for modeling the complex structures inherent in real-world bundles.
Building on these foundations, large language models (LLMs) have recently emerged as powerful tools for bundle recommendation, owing to their broad knowledge and strong reasoning capabilities (Liao et al., 2024; Zheng et al., 2024; Zhang et al., 2025). Existing efforts can be broadly grouped into two directions. The first focuses on improving item and bundle representations (Bayrak and Kaplan, 2025; Liu et al., 2025). For instance, (Bayrak and Kaplan, 2025) uses LLMs for offline data augmentation, enriching interaction data for embedding training. Similarly, Bundle-MLLM (Liu et al., 2025) unifies textual, visual, relational, and behavioral information into a tokenized input, allowing LLMs to form coherent multimodal representations through attention-based fusion. The second direction treats LLMs as generative engines for bundle construction (Liu et al., 2025; Sun et al., 2024c; Hu et al., 2025; Feng et al., 2025a, b). Representative examples include AICL (Sun et al., 2024c), which adapts LLMs with retrieval-augmented demonstrations and self-correction; Text2Bundle (Hu et al., 2025), which combines intent extraction with reinforcement learning for bundle generation; and RouteDK (Feng et al., 2025a), which employs knowledge distillation and mixture-of-experts to balance efficiency and knowledge consistency. Knowledge distillation has further been explored to reduce costs without degrading bundle generation quality (Feng et al., 2025b). These studies provide a more comprehensive and forward-looking perspective for LLMs in bundle recommendation.
Looking ahead, promising directions include positioning LLMs directly as bundle recommenders - either serving as semantic rankers over candidate bundles or evolving toward end-to-end generative recommenders such as OneRec (Deng et al., 2025). Nevertheless, several key challenges remain. First, scaling LLMs to handle the long and complex sequences involved in candidate bundle sets is critical. Second, stronger controllability over output formats is needed to ensure generated bundles meet system requirements. Third, reliability must be addressed through self-correction and feedback mechanisms to mitigate hallucinations during inference. Finally, more effective integration of multimodal content information (Sheng et al., 2025; Zheng et al., 2024)—including textual, visual, and behavioral data — is required to enable deeper personalization and stronger user-centric bundle recommendation.
@@ Acknowledgements.
This work is partially supported by NSFC, China (No.62276196), the research grant (RGPIN-2020-07157) from the Natural Science and Engineering Research Council (NSERC) of Canada and York Research Chairs (YRC) program. The authors also gratefully appreciate the anonymous reviewers and associate editor for their valuable comments and constructive suggestions that greatly helped to improve the quality of the paper.